15. For the process S, O So t g under continuous irradiation in different percentages of the quencher oxygen o) the Stern-Volmer linear fit was (I /In) 1.00 0.09596 x (%0,) Given k, 77 x 10 s s 1, calculate the fluorescence relaxation time when a short pulse is applied on the same complex with 50% O present. A) 9.74x10 3 s B) 5.65x10 s C) 1.84x106 s DO 9.74x10 s E) 5.85x10 7 sAnswer is A
Answer #1

There are two types of quenching mechanism. One is static and other is dynamic.

Oxygen is a dynamic quencher.

For dynamic quenching, lifetime varies

If0/If= t0/t, where is the initial lifetime t0 and is the t current lifetime.

Initial lifetime can be found from the fluorescence rate constant (t0= 1/kf) = 0.564 x 10-5s.

And the

If0/If = 1+ 0.09596 (% O2) at 50 % of O2

If0/If= 5.798

Using this equation

If0/If= t0/t

t= t0/(If0/If) = 0.564 * 10-5 / 5.798 = 0.097* 10-5 or 9.7* 10-3